OpenStudy (anonymous):

Find the exact value of sine of 4 times pi divided by 3.

Parth (parthkohli):

\(4\pi\) is just like 2 complete rounds, which is just no round if you look closely.

Parth (parthkohli):

What is \(\sin 0^{\circ}\)?

OpenStudy (anonymous):

0

Parth (parthkohli):

Yes, so what is \(0 \div 3\)?

OpenStudy (anonymous):

0 :P

Parth (parthkohli):

Ah you got it.

OpenStudy (anonymous):

but that isnt one of my options?

OpenStudy (anonymous):

the value is -\[\sqrt{3}\]/2

OpenStudy (anonymous):

negative square root 3 over 2

OpenStudy (anonymous):

or you can say -0.866 both are same u can check your options